OpenStudy (anonymous):

Solve the equation. Check for extraneous solutions. 2 l 7-7x l=2x+4

OpenStudy (anonymous):

A. x=3/2 or x=5/6 B. x=3/2 or x=5/8 C x=5/8 D. x=3/2

OpenStudy (anonymous):

Please HELP! =] Will FAN!

OpenStudy (anonymous):

@sweetburger

OpenStudy (muzzack):

I was thinking A

OpenStudy (anonymous):

I guessed B too!=]

OpenStudy (anonymous):

NEXT ONE,,,

OpenStudy (anonymous):

Is the inequality below sometimes, always, or never true? -2(2x + 9) > -4x + 9 A. always B. sometimes C. never

OpenStudy (the_fizicx99):

@zscdragon Refrain from giving out direct answers.

OpenStudy (anonymous):

Byyee Felicia!

OpenStudy (karatechopper):

Hold on Hold on Hold on, You can go to any college, university or whatever. But just know, that although you can use your smarts, you are using them in the wrong way. OpenStudy is HELP others. Prove your answer. Don't say it. Practice what you preach.